OpenStudy (anonymous):

Given , y=-3x^(2)+30x+7 complete the square and convert the equation to y=a(x − h)^(2) + k.

OpenStudy (anonymous):

i believe it is y=-3(x-5)^(2)+82 but im not sure i need help!

OpenStudy (anonymous):

lets see start with \[y=-3x^2+30x+7\] then \[y=-3(x^2-10x)+7\] then \[y=-3(x-5)^2+\text{some damned number}\]

OpenStudy (anonymous):

which we can find easily by replacing \(x\) by \(-5\) in the original expression to see what you get

Parth (parthkohli):

lol

OpenStudy (anonymous):

actually i was too busy being a wise guy replace \(x\) by \(5\) to see what you get

OpenStudy (anonymous):

you are correct

OpenStudy (anonymous):

\[-3\times (5)^2+30\times 5+7=82\]

OpenStudy (anonymous):

wait so, my first answer was right?